OpenStudy (anonymous):

Write the following equation in standard form: y + 9 = 4x – 8

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

Standard form is actually ax + by = c where a,b,c are all real numbers.

OpenStudy (anonymous):

so subtract 9 from both sides and subtract 4x. y - 4x = -17 or -4x + y = -17